Grade 11Algebra

prove 2+2=5

RAJ  RANJAN SINGH
15 Years agoGrade 11

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
ApprovedApproved Tutor Answer15 Years ago

Dear Raj

-20= -20

42-2*4*9/2= 52 - 2*5*9/2

42-2*4*9/2 + (9/2)2= 52 - 2*5*9/2 + (9/2)2

(4-9/2)2 = (5- 9/2)2

take square root

4-9/2 = 5-9/2

4=5

2+2=5

 

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

pranay -askiitians expert
ApprovedApproved Tutor Answer15 Years ago

Dear Raj

-20= -20

42-2*4*9/2= 52 - 2*5*9/2

42-2*4*9/2 + (9/2)2= 52 - 2*5*9/2 + (9/2)2

(4-9/2)2 = (5- 9/2)2

take square root

4-9/2 = 5-9/2

4=5

2+2=5

all the best